Maria’s preference over milk-shakes and fruit-smoothies is described by the utility function u(M, F) = M^3/4 + 3F, where M is the amount of milk-shakes and F the amount of fruit-smoothies she consumes every week. The price of milk-shakes is $4, and fruit-smoothies cost $8 per unit. Which of the following is true about Maria’s weekly consumption when she spends $16 every week on shakes and smoothies? (Assume that both goods are perfectly divisible)

a. Maria is consuming more fruit-smoothies than milk shakes

b. Maria would consume more milk shakes if she had more money to spend on the two goods.

c. Maria would consume less smoothies if their price increased to $9 per unit. (false) d. more than one of the above (false) e. none of the above (
Answer #1

UCM, F) = 134 +38 PM EH PF = 8 Budget = Blu At optimality MRS-Pmler 4M y = 2 mlynos M=(0.5) - 0.0625 UMA & F = 16 410.0625) +